A certain Carnot engine has Th = Tc + 35K, and an efficiency of 20.0%. What is Th?




A) 140 K B) 175 K C) 210 K D) 245 K


Please explain.
Answer #1

By Efficiency(E) = 1 - Tc/Th


=>20/100 = 1 - Tc/Th


=>Tc = 0.80Th -----------------(i)


=>Th = Tc + 35K (given) ------(ii)


=>By (i) & (ii) :-


=>0.20Th = 35K


=>Th =175 K
Answer #2

B) 175 K